OpenStudy (anonymous):

I need help understanding the system of equations if someone is willing to explain it to me. Please and thank you.

OpenStudy (anonymous):

What do you want to know about them? Basically a system of equations is just a set of 2 or more equations. If you want to solve a system of equations you want to find what values for the variables make all of the equations true. Graphically what you're trying to do is find where the graphs of all equations in the system intersect.

OpenStudy (anonymous):

@BangkokGarret well I need to cover the 3 graph, substitution, and elimination

OpenStudy (wolf1728):

By 3 graph do you mean solving for 3 unknowns?

OpenStudy (anonymous):

the 3 system @wolf1728

OpenStudy (wolf1728):

You mean solving for 3 unknowns correct?

OpenStudy (anonymous):

Yeah I believe so

OpenStudy (wolf1728):

I chose a rather simple set of equations: x + y + z = 6 x -y +z =2 -2x + y -z = -3 We need to eliminate 1 of the unknowns so we add all 3 equations and get y + z = 5 This allows us to get a value of z = 5 -y and we substitute this into the first equation x + y + 5 - y = 6 x = 1 Now you would need to find y and z.

OpenStudy (anonymous):

hm Okay o:

OpenStudy (wolf1728):

Actually, I made an easy example. For instance the solving for the 'x' unknown usually involves multiplying or dividing equations by certain numbers so that 2 unknowns can be isolated. This is a lucky coincidence that 2 unknowns can get eliminated in one step.

OpenStudy (anonymous):

Just starting out and my online tutors don't really explain very well so Im trying here so hopefully I can understand it. but im understanding it well.

OpenStudy (wolf1728):

To solve for y and z we have to use 2 equations and substitute the value of x=1, and make sure that another unknown gets eliminated. Luckily the first 2 equations work quite nicely. 1 + y + z = 6 1 -y + z =2 y + z = 5 -y + z = 1 Adding the 2 equations 2 z = 6 z = 3 From the first equation x +y +z = 6 so plugging in the values of x and z 1 + y +3 = 6 y = 2 Remember, solving for 3 unknowns is rarely this easy.

OpenStudy (wolf1728):

Maybe it's just me but it appears my first posting got shifted into the next to the last posting.

OpenStudy (anonymous):

Well could u explain to me Systems of Linear Equations? I'm suppose to learn that for my oral assessment

OpenStudy (wolf1728):

You need a system of equations when solving for more than one unknown. For 2 unknowns, you need 2 equations, for 3 unknowns you need 3 equations, etc.

OpenStudy (anonymous):

ah that makes sense

OpenStudy (anonymous):

Can you give me an example of a word problem?

OpenStudy (wolf1728):

I think a Google search would turn up thousands of examples of word problems. Anyway, I must get going. Perhaps someone else will be able to help you out. :-)

OpenStudy (anonymous):

Well thank you anyways.